OpenStudy (anonymous):

Will give a medal. Solve the system {7x+2y=5 {13x+14y=-1 using a matrix.

OpenStudy (anonymous):

I know the matrix will look like [ 7 2| 5] [13 14| -1] but don't know how to go forward with it.

OpenStudy (anonymous):

Step One: Line up the x, y and constant values. \[7x + 2y = 5 \]\[13x + 14y = -1\]Step Two: Write as equivalent matrices.\[\left[ 7x~~~~~2y \right]\]\[\left[ 13x~~~~14y \right]\]\[= \left[ 5\right]\]\[= \left[ -1 \right]\]

OpenStudy (anonymous):

But how do I solve using the matrix? As far as I know I'm supposed to multiply/divide and subtract each row to get to the end result. For example multiply row 7 2 5 by 2.

OpenStudy (anonymous):

However, I`m slightly confused on how to get 7 2 5 and 13 14 -1 equivalent to each other.

OpenStudy (anonymous):

Ok, thanks.